# GMAI-MMBench: A Comprehensive Multimodal Evaluation Benchmark Towards General Medical AI

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) are capable of handling diverse data types such as imaging, text, and physiological signals, and can be applied in various fields. In the medical field, LVLMs have a high potential to offer substantial assistance for diagnosis and treatment. Before that, it is crucial to develop benchmarks to evaluate LVLMs’ effectiveness in various medical applications. Current benchmarks are often built upon specific academic literature, mainly focusing on a single domain, and lacking varying perceptual granularities. Thus, they face specific challenges, including limited clinical relevance, incomplete evaluations, and insufficient guidance for interactive LVLMs. To address these limitations, we developed the GMAI-MMBench, the most comprehensive general medical AI benchmark with well-categorized data structure and multi-perceptual granularity to date. It is constructed from 285 datasets across 39 medical image modalities, 18 clinical-related tasks, 18 departments, and 4 perceptual granularities in a Visual Question Answering (VQA) format. Additionally, we implemented a lexical tree structure that allows users to customize evaluation tasks, accommodating various assessment needs and substantially supporting medical AI research and applications. We evaluated 50 LVLMs, and the results show that even the advanced GPT-4o only achieves an accuracy of 52%, indicating significant room for improvement. Moreover, we identified five key insufficiencies in current cutting-edge LVLMs that need to be addressed to advance the development of better medical applications. We believe that GMAI-MMBench will stimulate the community to build the next generation of LVLMs toward GMAI.
@@
OpenDataLab: https://opendatalab.com/GMAI/MMBench
Evaluation: https://github.com/open-compass/VLMEvalKit (Coming Soon)
@@ Introduction
| GMAI-MMBench | 39 | 26K | 18 | ✔ | I, B, M, C | 285 datasets from both public and hospital |
@@
First, GMAI-MMBench has diverse modalities and data sources because it is built upon 285 high-quality datasets collected across the world. These 285 datasets cover various medical image tasks, including 2D detection, 2D classification, and 2D/3D segmentation, to ensure the diversity of tasks. Using these foundational visual-based tasks has two advantages over using off-the-shelf image-text pair data. 1) It minimizes the risk of data leakage since the data in our benchmark are mostly image-label pairs rather than image-text pairs. The image-label pairs are not directly convertible to LVLMs training samples (usually image-text pairs), thus less likely to be used to train LVLMs; 2) It ensures high clinical relevance, as the images are sourced from hospitals and annotated by professional doctors. We then carefully selected approximately 26K cases with 39 different modalities to construct the GMAI-MMBench, thus meeting the modal diversity goal.
Second, GMAI-MMBench is a well-categorized medical benchmark that can comprehensively evaluate the pros and cons of various aspects of LVLMs, benefiting both model developers and users with specific needs. Specifically, we develop a categorization system, called lexical tree structure, which categorizes all cases into 18 clinical VQA tasks, 18 departments, 35 modalities, etc. The ‘clinical VQA tasks’ / ‘departments’ / ‘modalities’ are the lexicons that can be used to retrieve desired cases for evaluation. For instance, the oncology department can select cases related to oncology to evaluate LVLMs’ performance for oncology tasks, thus greatly enhancing flexibility and usability for specific demands.
@@
(1) GMAI-MMBench presents significant challenges in clinical practice. Even the best proprietary GPT-4o only achieves an accuracy of 52.24%, which demonstrates the deficiencies of cutting-edge LVLMs in tackling medical professional issues, thus they can hardly fulfill diverse clinical demands.
(2) Open-source LVLMs, such as MedDr and DeepSeek-VL-7B, achieve approximately 41% accuracy, making them very competitive compared to proprietary models. For instance, they surpass Claude3-Opus and Qwen-VL-Max and achieve comparable performance to Gemini 1.5 and GPT-4V. However, they still exhibit a clear performance disparity compared to the top-performing GPT-4o.
(3) Most medical-specific models have difficulty reaching the medium performance level (approximately 30% accuracy) achieved by general LVLMs, with the best performance achieved by MedDr.
@@
In summary, our contributions are three-fold. (a) We introduce a comprehensive benchmark, GMAI-MMBench, to evaluate existing LVLMs in clinical practice. GMAI-MMBench covers 39 modalities, 18 clinical VQA tasks, 18 departments, and 4 different perceptual granularity from 285 medical-related datasets, thereby offering a diverse range of modalities, tasks, and data sources. (b) GMAI-MMBench organizes each data point in lexical tree structures, with lexicons used to select desired data points to evaluate various aspects of LVLMs’ abilities. Thus, GMAI-MMBench facilitates customized evaluation to meet highly diverse demands in clinical practice. See Supplementary C.2. (c) We evaluate 44 representative general-purpose LVLMs, including both open-source and proprietary models, as well as 6 medical-specific LVLMs on GMAI-MMBench. The comprehensive evaluation reveals the pros and cons of different LVLMs from diverse perspectives, providing insights to improve these models to accommodate real-world clinical applications.
@@
We propose GMAI-MMBench, an innovative benchmark meticulously designed for the medical field, capable of providing comprehensive evaluations of LVLMs across various aspects of healthcare. (shown in the Figure2) We collect 285 datasets from public sources and hospitals, covering medical imaging tasks of detection, classification, and segmentation, to form the data fuel for establishing such a benchmark. The detailed datasets are listed in the supplementary. Based on the data foundation, we design a reliable pipeline to generate question-answering pairs and organize them from different perspectives with manual validation. Finally, we carefully select approximately 26K questions with varying levels of perceptual granularity from the manually validated cases to construct the final GMAI-MMBench.
@@
Dataset collection and standardization. As our aim is to build a large-scale benchmark for the comprehensive evaluation of LVLMs, the first and most important step is data collection. In contrast to benchmarks that directly use multimodal paired datasets, we source the datasets in two ways to minimize the data leakage problem and ensure the diversity and clinical property: First, we conduct thorough Internet searches to collect as many 2D/3D medical-related datasets as possible, retaining those that involve classification, detection, and segmentation tasks. Second, we collaborate with several hospitals that have agreed to share their ethically approved data. This process has enabled us to curate 285 datasets with highly qualified labels. Following data collection, we standardize both images and labels. For images, we adhere to the SA-Med2D-20M [257] protocol, transforming all 2D/3D medical images into 2D RGB images for further evaluation. For labels, we refer to the Medical Subject Headings (MeSH)111https://www.ncbi.nlm.nih.gov/mesh/1000048 to ensure every label is unique, clear, and free from conflict or ambiguity within each task. Specifically, we focus on three main situations: (1) expanding all abbreviations, such as changing “AMD” to “Age-related macular degeneration”; (2) unifying different expressions for the same target, such as standardizing both “lung nodule” and “pulmonary nodule” to “pulmonary nodule”; (3) merging labels with left and right distinctions, such as combining “left kidney” and “right kidney” into “kidney”, since our goal is to evaluate the abilities of understanding and reasoning rather than directional judgment.
@@ Results
Conclusion 1. Medical tasks are still challenging for all LVLMs: Our GMAI-MMBench provides a comprehensive multitask challenge, revealing that even the most advanced model, GPT-4o, is limited to an accuracy of around 52% (see Table 2 and Table 3). This does not meet the clinical requirement and indicates that all current LVLMs in the medical domain still require significant improvement.
@@
Insufficiency 1. Performance on different clinical VQA tasks needs improvement: Table 2 shows that the best-performing clinical VQA tasks are Disease Diagnosis (DD) and Nervous Tissue (NT), with models exceeding the random baseline by an average of over 10%. However, in clinical VQA tasks requiring complex reasoning such as Severity Grading (SG) and Image Quality Grading (IQG), most LVLMs face challenges, and most of them perform worse than the random baseline. Overall, despite the advanced models like GPT-4o and Gemini 1.5 significantly outperforming the random baseline, there remains a substantial gap between their performance and the requirements of real-world applications, indicating that all the models still need more specialized medical knowledge for training.
Insufficiency 2. The performance across different departments needs further balancing: In examining performance across different medical departments, as shown in Table 3, we found that the Cardiovascular Surgery (CS) and Neurosurgery (N) departments performed the best. In contrast, departments such as General Surgery (GS), Obstetrics and Gynecology (OG), and Urology (U) showed a need for improvement, as the performance of all models in these areas did not significantly exceed the random baseline compared to other departments. This indicates that current large models exhibit specialization biases, suggesting that future development of LVLMs aiming to achieve general medical AI should focus on balancing capabilities across all departments.
@@ Appendix B Dataset Details
In this section, we provide the detailed datasets used in GMAI-MMBench, including the name of the dataset or challenge, the number of sub-datasets in it, the modality, the dimension of data, the task type, and the number of cases. As shown in Table LABEL:tab:dataset_statistics, GMAI-MMBench is constructed from 285 datasets across 39 medical image modalities. These datasets are derived from the public (268) and several hospitals (17) that have agreed to share their ethically approved data.
@@
| 5K+ CT Images on Fractured Limbs [214] | 1 | CT | 2D | Cls | 30 |
@@
| Augemnted ocular diseases | 1 | Fundus Photography | 2D | Cls | 89 |
@@ C.1 Data Statistics
In this section, we present the comprehensive statistical information of GMAI-MMBench. Figure 6 offers a global view of the label distribution proportions for different clinical VQA tasks, departments, and perceptual granularities. The left pie chart (A) shows the distribution of clinical VQA tasks, with Disease Diagnosis (DD) being the most prevalent at 18.6%, followed by Surgical Instrument Recognition (SIR) at 9.5%, Surgical Workflow Recognition (SWR) at 8.6%, and Anatomical Structure Recognition (ASR) at 8.4%. The middle pie chart (B) depicts the distribution of cases across various departments, where Pulmonary Medicine (PM) has the highest proportion at 16.0%, followed by Hematology (H) at 10.3%, General Surgery (GS) at 10.0%, and Laboratory Medicine and Pathology (LMP) at 11.1%. The right pie chart (C) represents the distribution of perceptual granularities, with Image Level accounting for the largest share at 57.2%, followed by Mask Level at 17.1%, and Contour Level at 11.6%. Specifically, Table 5 provides the statistical details for different clinical VQA tasks, including their full terms, abbreviations, and the number of questions associated with each task. Table 6 presents the statistical information for different departments, including each department’s full term, abbreviation, and the number of questions contained within each department. Table 7 shows the statistical information for different granularity. In the detailed tables, the statistical information for multiple-choice questions is also included, specially, for multiple-choice questions, we count the frequency of choice appearances rather than the actual number of cases.
@@
| Anatomical Structure Recognition | ASR | 8 | 42 | 1733 | 1 | 18 | 7614 |
| Attribute Recognition | AR | 5 | 34 | 975 | 1 | 4 | 40 |
| Blood Vessels Recognition | BVR | 5 | 12 | 692 | - | - | - |
| Bone | B | 5 | 12 | 720 | - | - | - |
| Cell Recognition | CR | 4 | 13 | 676 | - | - | - |
@@
| Microorganism Recognition | MR | 2 | 26 | 809 | - | - | - |
| Muscle | M | 1 | 5 | 300 | - | - | - |
| Nervous Tissue | NT | 2 | 4 | 240 | - | - | - |
| Organ Recognition - Abdomen | OR-A | 4 | 10 | 600 | - | - | - |
| Organ Recognition - Head and Neck | OR-HN | 3 | 13 | 780 | - | - | - |
| Organ Recognition - Pelvic | OR-P | 5 | 6 | 360 | - | - | - |
| Organ Recognition - Thorax | OR-T | 7 | 15 | 900 | - | - | - |
| Severity Grading | SG | 6 | 65 | 1772 | - | - | - |
@@
| Surgical Instrument Recognition | SIR | 1 | 27 | 1406 | - | - | - |
@@
| Cardiovascular Surgery | CS | 8 | 9 | 510 | - | - | - |
| Dermatology | D | 1 | 30 | 1104 | - | - | - |
| Endocrinology | E | 3 | 7 | 300 | - | - | - |
| Gastroenterology and Hepatology | GH | 7 | 60 | 2374 | - | - | - |
| General Surgery | GS | 6 | 68 | 2685 | - | - | - |
| Hematology | H | 6 | 81 | 2472 | - | - | - |
| Infectious Diseases | ID | 2 | 6 | 180 | - | - | - |
| Laboratory Medicine and Pathology | LMP | 2 | 46 | 1469 | - | - | - |
| Nephrology and Hypertension | NH | 5 | 10 | 480 | - | - | - |
| Neurosurgery | N | 8 | 9 | 330 | - | - | - |
| Obstetrics and Gynecology | OG | 5 | 14 | 419 | - | - | - |
| Oncology (Medical) | OM | 20 | 51 | 1924 | - | - | - |
| Ophthalmology | O | 6 | 96 | 2680 | 2 | 11 | 218 |
| Orthopedic Surgery | OS | 8 | 55 | 2181 | - | - | - |
| Otolaryngology (ENT)/Head and Neck Surgery | ENT/HNS | 5 | 14 | 780 | 1 | 6 | 1015 |
| Pulmonary Medicine | PM | 2 | 55 | 2423 | 1 | 13 | 6844 |
| Sports Medicine | SM | 3 | 64 | 2069 | - | - | - |
| Urology | U | 8 | 33 | 1023 | - | - | - |
@@
| Mask Level | 36 | 177 | 11210 |
| Contour Level | 36 | 177 | 11210 |

### v4

@@ Abstract
Large Vision-Language Models (LVLMs) are capable of handling diverse data types such as imaging, text, and physiological signals, and can be applied in various fields. In the medical field, LVLMs have a high potential to offer substantial assistance for diagnosis and treatment. Before that, it is crucial to develop benchmarks to evaluate LVLMs’ effectiveness in various medical applications. Current benchmarks are often built upon specific academic literature, mainly focusing on a single domain, and lacking varying perceptual granularities. Thus, they face specific challenges, including limited clinical relevance, incomplete evaluations, and insufficient guidance for interactive LVLMs. To address these limitations, we developed the GMAI-MMBench, the most comprehensive general medical AI benchmark with well-categorized data structure and multi-perceptual granularity to date. It is constructed from 284 datasets across 38 medical image modalities, 18 clinical-related tasks, 18 departments, and 4 perceptual granularities in a Visual Question Answering (VQA) format. Additionally, we implemented a lexical tree structure that allows users to customize evaluation tasks, accommodating various assessment needs and substantially supporting medical AI research and applications. We evaluated 50 LVLMs, and the results show that even the advanced GPT-4o only achieves an accuracy of 53.96%, indicating significant room for improvement. Moreover, we identified five key insufficiencies in current cutting-edge LVLMs that need to be addressed to advance the development of better medical applications. We believe that GMAI-MMBench will stimulate the community to build the next generation of LVLMs toward GMAI.
@@ Introduction
| GMAI-MMBench | 38 | 26K | 18 | ✔ | I, B, M, C | 284 datasets from both public and hospital |
@@
First, GMAI-MMBench has diverse modalities and data sources because it is built upon 284 high-quality datasets collected across the world. These 284 datasets cover various medical image tasks, including 2D detection, 2D classification, and 2D/3D segmentation, to ensure the diversity of tasks. Using these foundational visual-based tasks has two advantages over using off-the-shelf image-text pair data. 1) It minimizes the risk of data leakage since the data in our benchmark are mostly image-label pairs rather than image-text pairs. The image-label pairs are not directly convertible to LVLMs training samples (usually image-text pairs), thus less likely to be used to train LVLMs; 2) It ensures high clinical relevance, as the images are sourced from hospitals and annotated by professional doctors. We then carefully selected approximately 26K cases with 38 different modalities to construct the GMAI-MMBench, thus meeting the modal diversity goal.
Second, GMAI-MMBench is a well-categorized medical benchmark that can comprehensively evaluate the pros and cons of various aspects of LVLMs, benefiting both model developers and users with specific needs. Specifically, we develop a categorization system, called lexical tree structure, which categorizes all cases into 18 clinical VQA tasks, 18 departments, 38 modalities, etc. The ‘clinical VQA tasks’ / ‘departments’ / ‘modalities’ are the lexicons that can be used to retrieve desired cases for evaluation. For instance, the oncology department can select cases related to oncology to evaluate LVLMs’ performance for oncology tasks, thus greatly enhancing flexibility and usability for specific demands.
@@
(1) GMAI-MMBench presents significant challenges in clinical practice. Even the best proprietary GPT-4o only achieves an accuracy of 53.96%, which demonstrates the deficiencies of cutting-edge LVLMs in tackling medical professional issues, thus they can hardly fulfill diverse clinical demands.
(2) Open-source LVLMs, such as MedDr and DeepSeek-VL-7B, achieve approximately 44% accuracy, making them very competitive compared to proprietary models. For instance, they surpass Claude3-Opus and Qwen-VL-Max and achieve comparable performance to Gemini 1.5 and GPT-4V. However, they still exhibit a clear performance disparity compared to the top-performing GPT-4o.
(3) Most medical-specific models have difficulty reaching a general performance level (approximately 30% accuracy) achieved by general LVLMs, except MedDr with 43.69% accuracy.
@@
In summary, our contributions are three-fold. (a) We introduce a comprehensive benchmark, GMAI-MMBench, to evaluate existing LVLMs in clinical practice. GMAI-MMBench covers 38 modalities, 18 clinical VQA tasks, 18 departments, and 4 different perceptual granularity from 284 medical-related datasets, thereby offering a diverse range of modalities, tasks, and data sources. (b) GMAI-MMBench organizes each data point in lexical tree structures, with lexicons used to select desired data points to evaluate various aspects of LVLMs’ abilities. Thus, GMAI-MMBench facilitates customized evaluation to meet highly diverse demands in clinical practice. See Supplementary C.2. (c) We evaluate 44 representative general-purpose LVLMs, including both open-source and proprietary models, as well as 6 medical-specific LVLMs on GMAI-MMBench. The comprehensive evaluation reveals the pros and cons of different LVLMs from diverse perspectives, providing insights to improve these models to accommodate real-world clinical applications.
@@
We propose GMAI-MMBench, an innovative benchmark meticulously designed for the medical field, capable of providing comprehensive evaluations of LVLMs across various aspects of healthcare. (shown in the Figure 2) We collect 284 datasets from public sources and hospitals, covering medical imaging tasks of detection, classification, and segmentation, to form the data fuel for establishing such a benchmark. The detailed datasets are listed in the supplementary. Based on the data foundation, we design a reliable pipeline to generate question-answering pairs and organize them from different perspectives with manual validation. Finally, we carefully select approximately 26K questions with varying levels of perceptual granularity from the manually validated cases to construct the final GMAI-MMBench.
@@
Dataset collection and standardization. As our aim is to build a large-scale benchmark for the comprehensive evaluation of LVLMs, the first and most important step is data collection. In contrast to benchmarks that directly use multimodal paired datasets, we source the datasets in two ways to minimize the data leakage problem and ensure the diversity and clinical property: First, we conduct thorough Internet searches to collect as many 2D/3D medical-related datasets as possible, retaining those that involve classification, detection, and segmentation tasks. Second, we collaborate with several hospitals that have agreed to share their ethically approved data. This process has enabled us to curate 284 datasets with highly qualified labels. Following data collection, we standardize both images and labels. For images, we adhere to the SA-Med2D-20M [258] protocol, transforming all 2D/3D medical images into 2D RGB images for further evaluation. For labels, we refer to the Medical Subject Headings (MeSH)111https://www.ncbi.nlm.nih.gov/mesh/1000048 to ensure every label is unique, clear, and free from conflict or ambiguity within each task. Specifically, we focus on three main situations: (1) expanding all abbreviations, such as changing “AMD” to “Age-related macular degeneration”; (2) unifying different expressions for the same target, such as standardizing both “lung nodule” and “pulmonary nodule” to “pulmonary nodule”; (3) merging labels with left and right distinctions, such as combining “left kidney” and “right kidney” into “kidney”, since our goal is to evaluate the abilities of understanding and reasoning rather than directional judgment.
@@ Results
Conclusion 1. Medical tasks are still challenging for all LVLMs: Our GMAI-MMBench provides a comprehensive multitask challenge, revealing that even the most advanced model, GPT-4o, is limited to an accuracy of around 54% (see Table 2 and Table 3). This does not meet the clinical requirement and indicates that all current LVLMs in the medical domain still require significant improvement.
@@
Insufficiency 1. Performance on different clinical VQA tasks needs improvement: Table 2 shows that the best-performing clinical VQA tasks are Disease Diagnosis (DD) and Nervous Tissue (NT), with models exceeding the random baseline by an average of over 10%. However, in clinical VQA tasks such as Severity Grading (SG) and Attribute Recognition (AR), most LVLMs face challenges, and most of them perform worse than the random baseline. Overall, despite the advanced models like GPT-4o and Gemini 1.5 significantly outperforming the random baseline, there remains a substantial gap between their performance and the requirements of real-world applications, indicating that all the models still need more specialized medical knowledge for training.
Insufficiency 2. The performance across different departments needs further balancing: In examining performance across different medical departments, as shown in Table 3, we found that the Infectious Diseases (ID) and Neurosurgery (N) departments performed the best. In contrast, departments such as General Surgery (GS) and Obstetrics and Gynecology (OG) showed a need for improvement, as the performance of all models in these areas did not significantly exceed the random baseline compared to other departments. This indicates that current large models exhibit specialization biases, suggesting that future development of LVLMs aiming to achieve general medical AI should focus on balancing capabilities across all departments.
@@ Appendix B Dataset Details
In this section, we provide the detailed datasets used in GMAI-MMBench, including the name of the dataset or challenge, the number of sub-datasets in it, the modality, the dimension of data, the task type, and the number of cases. As shown in Table LABEL:tab:dataset_statistics, GMAI-MMBench is constructed from 284 datasets across 38 medical image modalities. These datasets are derived from the public (268) and several hospitals (16) that have agreed to share their ethically approved data.
@@
| 5K+ CT Images on Fractured Limbs [215] | 1 | CT | 2D | Cls | 60 |
@@
| Augemnted ocular diseases | 1 | Fundus Photography | 2D | Cls | 97 |
@@ C.1 Data Statistics
In this section, we present the comprehensive statistical information of GMAI-MMBench. Figure 6 offers a global view of the label distribution proportions for different clinical VQA tasks, departments, and perceptual granularities. The left pie chart (A) shows the distribution of clinical VQA tasks, with Disease Diagnosis (DD) being the most prevalent at 51.6%, followed by Severity Grading (SG) at 9.1%, Counting (C) at 5.4%, and Organ Recognition – Abdomen (OR-A) at 4.0%. The middle pie chart (B) depicts the distribution of cases across various departments, where Ophthalmolog (O) has the highest proportion at 11.3%, followed by Hematology (H) at 10.7%, General Surgery (GS) at 10.2%, and Urolog (U) at 9.7%. The right pie chart (C) represents the distribution of perceptual granularities, with Image Level accounting for the largest share at 49.2%, followed by Mask Level at 22.0%, and Contour Level at 22.0%. Specifically, Table 5 provides the statistical details for different clinical VQA tasks, including their full terms, abbreviations, and the number of questions associated with each task. Table 6 presents the statistical information for different departments, including each department’s full term, abbreviation, and the number of questions contained within each department. Table 7 shows the statistical information for different granularity. In the detailed tables, the statistical information for multiple-choice questions is also included, specially, for multiple-choice questions, we count the frequency of choice appearances rather than the actual number of cases.
@@
| Attribute Recognition | AR | 5 | 26 | 780 | 1 | 4 | 40 |
| Blood Vessels Recognition | BVR | 7 | 15 | 436 | - | - | - |
| Bone | B | 6 | 22 | 655 | - | - | - |
| Cell Recognition | CR | 4 | 13 | 383 | 1 | 18 | 7614 |
@@
| Microorganism Recognition | MR | 3 | 26 | 779 | - | - | - |
| Muscle | M | 1 | 5 | 150 | - | - | - |
| Nervous Tissue | NT | 2 | 4 | 120 | - | - | - |
| Organ Recognition - Abdomen | OR-A | 7 | 28 | 838 | - | - | - |
| Organ Recognition - Head and Neck | OR-HN | 5 | 16 | 480 | - | - | - |
| Organ Recognition - Pelvic | OR-P | 6 | 9 | 270 | - | - | - |
| Organ Recognition - Thorax | OR-T | 9 | 17 | 510 | - | - | - |
| Severity Grading | SG | 5 | 64 | 1678 | - | - | - |
@@
| Surgical Instrument Recognition | SIR | 1 | 27 | 790 | - | - | - |
@@
| Cardiovascular Surgery | CS | 9 | 9 | 270 | 1 | 1 | 424 |
| Dermatology | D | 1 | 30 | 894 | - | - | - |
| Endocrinology | E | 3 | 7 | 210 | - | - | - |
| Gastroenterology and Hepatology | GH | 7 | 60 | 1774 | - | - | - |
| General Surgery | GS | 6 | 68 | 2009 | - | - | - |
| Hematology | H | 6 | 80 | 2112 | - | - | - |
| Infectious Diseases | ID | 2 | 7 | 180 | - | - | - |
| Laboratory Medicine and Pathology | LMP | 2 | 45 | 1259 | 1 | 18 | 7614 |
| Nephrology and Hypertension | NH | 4 | 9 | 270 | - | - | - |
| Neurosurgery | N | 8 | 9 | 270 | - | - | - |
| None (Attributes that do not belong to any department) | N/A | 2 | 15 | 450 | - | - | - |
| Obstetrics and Gynecology | OG | 5 | 14 | 389 | - | - | - |
| Oncology (Medical) | OM | 20 | 51 | 1399 | - | - | - |
| Ophthalmology | O | 6 | 97 | 2232 | 2 | 11 | 218 |
| Orthopedic Surgery | OS | 8 | 54 | 1611 | - | - | - |
| Otolaryngology (ENT)/Head and Neck Surgery | ENT/HNS | 5 | 14 | 420 | 1 | 6 | 1015 |
| Pulmonary Medicine | PM | 2 | 55 | 1643 | 1 | 12 | 6420 |
| Sports Medicine | SM | 3 | 64 | 1919 | - | - | - |
| Urology | U | 8 | 33 | 933 | - | - | - |
@@
| Mask Level | 36 | 188 | 5587 |
| Contour Level | 36 | 188 | 5587 |